# The Geometric Inductive Bias of Grokking: Bypassing Phase Transitions via Architectural Topology

### 05 Mar 2026 v1

@@ Abstract
Mechanistic interpretability has largely focused on post-hoc analysis of trained networks. In this work, we instead adopt an interventional approach: we test mechanistic hypotheses a priori by modifying architectural topology and observing the resulting training dynamics. We study grokking—the delayed generalization phenomenon observed when Transformers are trained on cyclic modular addition ($\mathbb{Z}_{p}$)—and investigate whether specific architectural degrees of freedom contribute to the prolonged memorization phase.
We identify two independent structural factors present in standard Transformers: unbounded representational magnitude and data-dependent attention routing. First, we introduce a fully bounded spherical topology that enforces $L_{2}$ normalization throughout the residual stream and normalizes the unembedding matrix with a fixed temperature scale. This constraint removes magnitude-based degrees of freedom and reduces grokking onset time by more than $20\times$ without weight decay.
Second, we introduce a Uniform Attention Ablation that overrides data-dependent query–key routing with a fixed uniform distribution, reducing the attention layer to a Continuous Bag-of-Words (CBOW) aggregator. Despite eliminating adaptive routing, LayerNorm models with uniform attention achieve 100% generalization across all seeds and bypass the grokking delay.
To evaluate whether this acceleration reflects task-specific geometric alignment rather than a generic optimization stabilizer, we employ non-commutative $S_{5}$ permutation composition as a negative control. Enforcing the same spherical constraints on $S_{5}$ does not accelerate generalization, suggesting that eliminating the memorization phase depends strongly on aligning architectural priors with the intrinsic symmetries of the task. Together, these findings provide interventional evidence that architectural degrees of freedom substantially influence the presence and duration of grokking, suggesting a predictive architectural perspective on training dynamics.
@@ 1 Introduction
In standard Transformers, information can be encoded in both the direction and magnitude of residual stream vectors. Prior work has suggested that unconstrained magnitude growth can influence training dynamics in grokking settings (Prieto et al., 2025). To examine the role of this degree of freedom, we introduce a Fully Bounded Spherical Topology that enforces strict $L_{2}$ normalization throughout the residual stream and applies a normalized unembedding matrix with a fixed temperature scale. This intervention removes the model’s ability to encode information in vector norm and constrains logits to a bounded cosine geometry. Empirically, this modification substantially accelerates generalization, reducing grokking onset from $\sim$54,160 epochs in standard LayerNorm baselines to $\sim$2,100 epochs without weight decay.
@@ 3.2 The Spherical Residual Stream and Fully Bounded Topology (Intervention A)
where $\epsilon=10^{-8}$ ensures numerical stability. This operator is applied to the residual stream prior to each sub-layer and immediately after each residual addition. A single transformer block is thus modified as:
@@
Fully Bounded Topology.
While the spherical residual stream provides the necessary geometry, leaving the output layer unconstrained allows the final linear unembedding matrix ($W_{\text{unembed}}$) to scale arbitrarily. Training under cross-entropy loss can drive increasing logit magnitudes to artificially lower the loss. This behavior—known as Naïve Loss Minimization—leads to numerical instability and Softmax Collapse, which can severely delay or crash generalization Prieto et al. (2025).
To stabilize the geometry and prevent Softmax Collapse, we additionally apply $\Pi_{S}$ to the unembedding weights and compute logits via scaled cosine similarity:
@@ 4 Empirical Results
In this section, we evaluate the hypothesis that excess representational degrees of freedom—specifically, residual magnitude flexibility and data-dependent attention routing—contribute to delayed generalization in modular arithmetic tasks. We evaluate our two structural interventions across standard baselines (LayerNorm, RMSNorm) and our bounded configurations (Spherical Norm with $\lambda=1.0$, and Fully Bounded with $\lambda=0.0$). All reported metrics are aggregated across 10 independent random seeds per configuration.
@@ 4.1 Intervention A: Accelerating Grokking via Topological Constraint
Spherical Norm ($\lambda=1.0$) 2,480 464 1,800 3,200 0 / 10 100%
Fully Bounded ($\lambda=0.0$) 2,100 316 1,800 2,600 0 / 10 100%
@@
Spherical Norm ($\lambda=1.0$) 820 199 600 1,200 0 / 10 100%
Fully Bounded ($\lambda=0.0$) 700 194 400 1,000 0 / 10 100%
In contrast to the baselines, the spherical configurations exhibit substantially earlier generalization. At $10^{-4}$, the Fully Bounded topology reaches 100% test accuracy in a mean of 2,100 epochs—over an order of magnitude faster than the statistical normalization baselines. At a higher learning rate ($6\times 10^{-4}$), baselines require approximately 7,500 epochs on average, while the Fully Bounded configuration generalizes in roughly 700 epochs.
@@
As theorized in Section 3.2, standard unconstrained models rely on explicit regularization to counteract Naïve Loss Minimization and prevent Softmax Collapse Prieto et al. (2025). When the spherical constraint was applied exclusively to the residual stream while retaining standard weight decay ($\lambda=1.0$), the model exhibited pronounced optimization instability. In this configuration, logit confidence must be achieved through scaling in the unconstrained unembedding layer, while weight decay simultaneously penalizes such growth. Empirically, this interaction coincided with large gradient spikes and variance across seeds.
In contrast, the fully bounded spherical topology—employing a normalized unembedding layer and zero weight decay ($\lambda=0.0$)—eliminates unbounded logit scaling as a degree of freedom. Under this configuration, gradient norms remained smooth, and the model reached 100% generalization rapidly. This indicates that under the fully bounded topology, stable optimization can emerge without reliance on weight decay regularization.
@@ 4.2 Intervention B: Bypassing Grokking via Uniform Attention Ablation
Spherical Norm (WD=1.0) 97.65% 100.00% 6 / 10
Fully Bounded Sphere (WD=0.0) 100.00% 100.00% 10 / 10
As shown in Table 2, the standard LayerNorm baseline equipped with uniform attention achieves 100% peak test accuracy across all 10 independent seeds, entirely bypassing the prolonged grokking plateau. The Fully Bounded spherical topology likewise maintains 100% success under uniform attention.
Beyond peak accuracy, examining the training dynamics under uniform attention reveals the impact of topological constraints on optimization stability. As illustrated in Figure 2, both the standard LayerNorm baseline (Figure 2a) and the Fully Bounded topology (Figure 2c) exhibit an immediate, concurrent rise in training and test accuracy, bypassing the extended memorization plateau observed in other settings. In contrast, when the spherical constraint is applied while retaining weight decay (Figure 2b), the optimization trajectory becomes markedly unstable. The tension between the model increasing logit magnitudes to reduce cross-entropy loss and weight decay penalizing that growth appears to introduce substantial optimization friction, resulting in delayed generalization.
@@
(b) Spherical Norm ($\lambda=1.0$)
(c) Fully Bounded ($\lambda=0.0$)
Figure 2: Training dynamics under Uniform Attention Ablation (Seed 2). (a) The LayerNorm baseline immediately generalizes when data-dependent routing is removed, bypassing grokking. (b) Imposing a spherical constraint while retaining weight decay introduces optimization instability, delaying generalization. (c) The Fully Bounded topology (zero weight decay) stabilizes the geometry, resulting in immediate and flawless generalization.
@@ 4.3 Spectral Verification of the Fourier Circuit
Optimization Friction (Sphere, $\lambda=1.0$).
When spherical normalization is applied to the residual stream while retaining an unconstrained unembedding and standard weight decay ($\lambda=1.0$), circuit coherence degrades substantially. The maximum FVE does not exceed $\sim 29\%$, with several dominant frequencies explaining less than 10% of activation variance. This fragmentation is consistent with the optimization instability arising from the interaction between loss-driven logit scaling and weight decay regularization.
Fully Bounded Manifold (Sphere, $\lambda=0.0$).
In contrast, the fully bounded topology (normalized unembedding, $\lambda=0.0$) exhibits strong spectral alignment. Dominant frequencies explain a substantial fraction of activation variance (e.g., FVE U: 62.55%, V: 48.43%). By removing magnitude scaling from both the residual stream and the unembedding layer, optimization converges rapidly to a geometrically coherent Fourier representation.
Bounding the residual stream alone does not reliably produce coherent Fourier structure; stronger spectral alignment is observed when magnitude scaling is eliminated throughout the full pipeline.
@@ 4.4 Task-Specific Alignment: Negative Control on $S_{5}$
Spherical Norm ($\lambda=1.0$) Failed 10 / 10
Fully Bounded ($\lambda=0.0$) Failed 10 / 10
@@
This differential outcome provides evidence that the Fully Bounded topology functions as a task-specific geometric inductive bias rather than a generic optimization accelerator. If the spherical constraint merely improved optimization dynamics, comparable acceleration would be expected across tasks. Instead, we observe acceleration on modular addition but consistent failure on $S_{5}$ under the same training regime.
@@ 5.3 The Bitter Lesson and Modality-Specific Debugging
Transformers are increasingly deployed beyond text, including in time-series forecasting, reinforcement learning, and structured decision-making tasks. In such settings, recent architectural interventions provide suggestive empirical support for symmetry-aligned design. For example, Gillman et al. (2025) demonstrate that replacing a standard linear classification head with a topologically constrained “Fourier head” significantly improves performance on continuous decision-making tasks. Similarly, Sun et al. (2025) show that explicitly incorporating periodic and Fourier-based group attention mechanisms in the PENGUIN architecture yields strong results in long-horizon time-series forecasting.
@@ A.2 Task 1: Modular Addition ($\mathbb{Z}_{113}$)
  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Fully Bounded Sphere $\lambda=0.0$)
@@ A.3 Task 2: Permutation Composition ($S_{5}$)
  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Fully Bounded Sphere $\lambda=0.0$)

### 10 Mar 2026 v2

@@ Abstract
Mechanistic interpretability typically relies on post-hoc analysis of trained networks. We instead adopt an interventional approach: testing hypotheses a priori by modifying architectural topology to observe training dynamics. We study grokking—delayed generalization in Transformers trained on cyclic modular addition ($\mathbb{Z}_{p}$)—investigating if specific architectural degrees of freedom prolong the memorization phase.
We identify two independent structural factors in standard Transformers: unbounded representational magnitude and data-dependent attention routing. First, we introduce a fully bounded spherical topology enforcing $L_{2}$ normalization throughout the residual stream and an unembedding matrix with a fixed temperature scale. This removes magnitude-based degrees of freedom and qualitatively alters the training dynamics: models generalize immediately, bypassing the prolonged memorization phase characteristic of grokking in standard Transformer baselines, both with and without weight decay. Second, a Uniform Attention Ablation overrides data-dependent query-key routing with a uniform distribution, reducing the attention layer to a Continuous Bag-of-Words (CBOW) aggregator. Despite removing adaptive routing, these models achieve $100\%$ generalization across all seeds and bypass the grokking delay entirely.
To evaluate whether this acceleration is a task-specific geometric alignment rather than a generic optimization stabilizer, we use non-commutative $S_{5}$ permutation composition as a negative control. Enforcing spherical constraints on $S_{5}$ does not accelerate generalization. This suggests eliminating the memorization phase depends strongly on aligning architectural priors with the task’s intrinsic symmetries. Together, these findings provide interventional evidence that architectural degrees of freedom substantially influence grokking, suggesting a predictive structural perspective on training dynamics.
@@ 1 Introduction
In standard Transformers, information can be encoded in both the direction and magnitude of residual stream vectors. Prior work has suggested that unconstrained magnitude growth can influence training dynamics in grokking settings (Prieto et al., 2025). To examine the role of this degree of freedom, we introduce a Fully Bounded Spherical Topology that enforces strict $L_{2}$ normalization throughout the residual stream and applies a normalized unembedding matrix with a fixed temperature scale. This intervention removes the model’s ability to encode information in vector norm and constrains logits to a bounded cosine geometry. Empirically, this modification substantially accelerates generalization, reducing grokking onset from $\sim$54,160 epochs in standard LayerNorm baselines to $\sim$2,480 epochs without weight decay.
@@ 2.3 Fourier Features and Structural Continuity
Accelerating Grokking via Optimization.
Complementary to geometric and representational analyses, Lee et al. (2024) demonstrated that grokking can be accelerated at the optimizer level by amplifying slow-varying gradient components via low-pass filtering. Their GROKFAST algorithm achieves up to $50\times$ faster generalization by boosting the low-frequency signal responsible for the delayed phase transition. However, this approach operates within the standard memorize-then-generalize paradigm: the model still undergoes a memorization phase before transitioning to generalization. In contrast, our architectural interventions eliminate the memorization phase entirely—bounded spherical models generalize from initialization without ever entering the prolonged overfitting plateau.
@@ 3.2 The Spherical Residual Stream and Fully Bounded Topology (Intervention A)
where $\epsilon=10^{-12}$ ensures numerical stability. This operator is applied to the residual stream prior to each sub-layer and immediately after each residual addition. A single transformer block is thus modified as:
@@
Bounded Output Topology.
The spherical residual stream alone is insufficient: if the final unembedding matrix ($W_{\text{unembed}}$) remains unconstrained, training under cross-entropy loss can drive increasing logit magnitudes to artificially lower the loss. This behavior—known as Naïve Loss Minimization—leads to numerical instability and Softmax Collapse, which can severely delay or crash generalization Prieto et al. (2025). Empirically, we found that all spherical models required unembedding normalization to train stably.
To stabilize the geometry and prevent Softmax Collapse, we apply $\Pi_{S}$ to the unembedding weights and compute logits via scaled cosine similarity:
@@
Fourier Initialization.
In addition to the standard random initialization, we evaluate a variant in which the first 10 dimensions of the token embedding matrix are deterministically initialized with cosine and sine values at five key frequencies, following the Fourier structure identified in prior mechanistic analyses (Nanda et al., 2023). This provides the model with an explicit periodic prior from initialization. We refer to this variant as Bounded Sphere + Fourier Init.
@@ 4 Empirical Results
In this section, we evaluate the hypothesis that excess representational degrees of freedom—specifically, residual magnitude flexibility and data-dependent attention routing—contribute to delayed generalization in modular arithmetic tasks. We evaluate our two structural interventions across standard baselines (LayerNorm, RMSNorm) and our bounded spherical topology tested with weight decay ($\lambda=1.0$) and without ($\lambda=0.0$). All reported metrics are aggregated across 10 independent random seeds per configuration.
@@ 4.1 Intervention A: Accelerating Grokking via Topological Constraint
Bounded Sphere ($\lambda=1.0$) 2,400 353 1,800 3,000 0 / 10 100%
Bounded Sphere ($\lambda=0.0$) 2,480 464 1,800 3,200 0 / 10 100%
Bounded Sphere + Fourier Init ($\lambda=1.0$) 2,120 368 1,600 2,600 0 / 10 100%
Bounded Sphere + Fourier Init ($\lambda=0.0$) 2,100 316 1,800 2,600 0 / 10 100%
@@
Bounded Sphere ($\lambda=1.0$) 1,560 853 600 3,000 0 / 10 100%
Bounded Sphere ($\lambda=0.0$) 820 199 600 1,200 0 / 10 100%
Bounded Sphere + Fourier Init ($\lambda=1.0$) 800 163 600 1,000 0 / 10 100%
Bounded Sphere + Fourier Init ($\lambda=0.0$) 711 203 400 1,000 0 / 10 100%
In contrast to the baselines, the spherical configurations exhibit substantially earlier generalization. At $10^{-4}$, the Bounded Sphere ($\lambda=0.0$) reaches 100% test accuracy in a mean of 2,480 epochs—over an order of magnitude faster than the statistical normalization baselines. At a higher learning rate ($6\times 10^{-4}$), baselines require approximately 7,500 epochs on average, while the Bounded Sphere ($\lambda=0.0$) generalizes in roughly 820 epochs.
@@
As theorized in Section 3.2, standard unconstrained models rely on explicit regularization to counteract Naïve Loss Minimization and prevent Softmax Collapse Prieto et al. (2025). All spherical configurations employ the same bounded architecture: $L_{2}$-normalized residual stream, normalized unembedding, and fixed temperature scaling. The only variable is weight decay. At the lower learning rate ($1\times 10^{-4}$), weight decay has minimal effect—both $\lambda=1.0$ and $\lambda=0.0$ converge in $\sim$2,400–2,480 epochs. At the higher learning rate ($6\times 10^{-4}$), however, weight decay introduces optimization friction: the $\lambda=1.0$ configuration exhibits higher variance across seeds (std 853 vs. 199) and slower mean convergence (1,560 vs. 820 epochs). This suggests that under the bounded topology, weight decay penalizes weight magnitudes that are already geometrically constrained, introducing unnecessary regularization pressure at aggressive learning rates.
Crucially, bounding the unembedding alone does not eliminate delayed generalization. When an $L_{2}$-normalized unembedding is applied to the LayerNorm baseline, the model still exhibits a prolonged memorization phase. This occurs because LayerNorm preserves internal radial freedom in the residual stream, allowing the network to increase residual magnitudes in order to reduce cross-entropy loss through Naïve Loss Minimization Prieto et al. (2025). While recent work proposes numerical stabilizations (e.g., StableMax) that allow models to remain numerically stable during this unconstrained scaling phase Prieto et al. (2025), our bounded spherical topology instead removes this internal magnitude degree of freedom at the architectural level. By structurally constraining the residual stream itself, the architecture bypasses the prolonged memorization phase observed in baseline models and exhibits immediate generalization.
@@ 4.2 Intervention B: Bypassing Grokking via Uniform Attention Ablation
Bounded Sphere ($\lambda=1.0$) 97.65% 100.00% 6 / 10
Bounded Sphere ($\lambda=0.0$) 100.00% 100.00% 10 / 10
As shown in Table 2, the standard LayerNorm baseline equipped with uniform attention achieves 100% peak test accuracy across all 10 independent seeds, entirely bypassing the prolonged grokking plateau. The Bounded Sphere ($\lambda=0.0$) likewise maintains 100% success under uniform attention.
Beyond peak accuracy, examining the training dynamics under uniform attention reveals the impact of topological constraints on optimization stability. As illustrated in Figure 2, both the standard LayerNorm baseline (Figure 2a) and the Bounded Sphere ($\lambda=0.0$) (Figure 2c) exhibit an immediate, concurrent rise in training and test accuracy, bypassing the extended memorization plateau observed in other settings. In contrast, when the spherical constraint is applied while retaining weight decay (Figure 2b), the optimization trajectory becomes markedly unstable. The tension between the model increasing logit magnitudes to reduce cross-entropy loss and weight decay penalizing that growth appears to introduce substantial optimization friction, resulting in delayed generalization.
@@
(b) Bounded Sphere ($\lambda=1.0$)
(c) Bounded Sphere ($\lambda=0.0$)
Figure 2: Training dynamics under Uniform Attention Ablation (Seed 2). (a) The LayerNorm baseline immediately generalizes when data-dependent routing is removed, bypassing grokking. (b) Imposing a spherical constraint while retaining weight decay introduces optimization instability, delaying generalization. (c) The Bounded Sphere without weight decay ($\lambda=0.0$) stabilizes the geometry, resulting in immediate and flawless generalization.
@@ 4.3 Spectral Verification of the Fourier Circuit
Optimization Friction (Bounded Sphere, $\lambda=1.0$).
When the bounded spherical topology is trained with standard weight decay ($\lambda=1.0$), circuit coherence degrades substantially. The maximum FVE does not exceed $\sim 29\%$, with several dominant frequencies explaining less than 10% of activation variance. This fragmentation is consistent with the optimization instability arising from the interaction between loss-driven logit scaling and weight decay regularization.
Bounded Sphere ($\lambda=0.0$).
In contrast, the same bounded topology trained without weight decay ($\lambda=0.0$) exhibits strong spectral alignment. Dominant frequencies explain a substantial fraction of activation variance (e.g., FVE U: 62.55%, V: 48.43%). By removing magnitude scaling from both the residual stream and the unembedding layer, optimization converges rapidly to a geometrically coherent Fourier representation.
@@ 4.4 Task-Specific Alignment: Negative Control on $S_{5}$
Bounded Sphere ($\lambda=1.0$) Failed 10 / 10
Bounded Sphere ($\lambda=0.0$) Failed 10 / 10
@@
This differential outcome provides evidence that the bounded spherical topology functions as a task-specific geometric inductive bias rather than a generic optimization accelerator. If the spherical constraint merely improved optimization dynamics, comparable acceleration would be expected across tasks. Instead, we observe acceleration on modular addition but consistent failure on $S_{5}$ under the same training regime.
@@ 5.3 The Bitter Lesson and Modality-Specific Debugging
Transformers are increasingly deployed beyond text, including in time-series forecasting, reinforcement learning, and structured decision-making tasks. In such settings, recent architectural interventions provide suggestive empirical support for symmetry-aligned design. For example, Gillman et al. (2025) demonstrate that replacing a standard linear classification head with a topologically constrained “Fourier head” significantly improves performance on continuous decision-making tasks. Similarly, Sun et al. (2025) show that explicitly incorporating a modulo-based periodic relative attention bias in the PENGUIN architecture yields strong results in long-horizon time-series forecasting by structurally aligning the attention mechanism with the data’s intrinsic cyclic symmetry.
@@ 5.5 Conclusion
Acknowledgements
Large language models were used in the preparation of this manuscript for limited assistance with language polishing, LaTeX formatting of tables and figures, and routine programming tasks. All experimental design, implementation, analysis, and scientific conclusions were developed and verified by the author, who takes full responsibility for the content of this work.
@@ A.2 Task 1: Modular Addition ($\mathbb{Z}_{113}$)
  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Bounded Sphere $\lambda=0.0$)
@@ A.3 Task 2: Permutation Composition ($S_{5}$)
  • Weight Decay: $1.0$ (Baselines & Sphere $\lambda=1.0$) / $0.0$ (Bounded Sphere $\lambda=0.0$)